OpenStudy (anonymous):

The following geometric sequences represent the populations of two bacterial cultures after 0 hours, 1 hour, 2 hours, and so on. Culture A starts with more bacteria, but culture B has a ratio of increase that is twice as large. Which culture will have the greatest population after 10 hours? Culture A: 1000, 1500, 2250, 3375, ... Culture B: 1, 3, 9, 27, ... A. Culture A B. Culture B

OpenStudy (anonymous):

help

OpenStudy (kash_thesmartguy):

First we need to understand the sequences. According to my calculations- #1 doubles the value it adds. #2 (*3) the value

OpenStudy (anonymous):

ok

OpenStudy (welshfella):

the population after 10 hours will be a1 r^9 where a = starting population and r = common ratio

OpenStudy (welshfella):

*a1

OpenStudy (anonymous):

so a is 2hours

OpenStudy (anonymous):

so its b

OpenStudy (welshfella):

the common ratio for A is 1500/1000 = 1.5 and for b its 3/1 = 3

OpenStudy (welshfella):

after 10years A = 1000* 1.5^9 and B = 1* 3^9

OpenStudy (anonymous):

so its a

OpenStudy (anonymous):

right????

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

thx

OpenStudy (welshfella):

A = 38443 and B = 19683

OpenStudy (welshfella):

remember a GP has term a1, a1 r, a1r^2, a1r^3....

OpenStudy (anonymous):

kk